[15pts] An FeCls solution required 301 s of a 2.74 A current to pass through in order to plate 0.150 g of iron out of solution. What was the plating efficiency? 6.
Answer #1

6)

Electrolysis equation is:

Fe3+ + 3e- ------> Fe

1 mol of Fe requires 3 mol of electron

1 mol of electron = 96485 C

So,1 mol of Fe requires 289455 C

let us calculate the charge passed:

time, t = 3.01*10^2s

Q = I*t

= 2.74A * 3.01*10^2s

= 8.247*10^2 C

mol of Fe plated = 8.247*10^2/289455 = 2.849*10^-3 mol

Molar mass of Fe = 55.85 g/mol

mass of Fe = number of mol * molar mass

= 2.849*10^-3 * 55.85

= 0.159 g

This should be theoretical mass

But actual mass = 0.150 g

So,

efficiency = actual mass * 100 / theoretical mass

= 0.150 * 100 / 0.159

= 94.3 %

Answer: 94.3 %